# Computational Economics in Large Language Models: Exploring Model Behavior and Incentive Design under Resource Constraints

### v2

@@ 5.3.2 Ablation Study: The Source of Savings
To disentangle the sources of efficiency gains, we applied the incentive penalty to only the attention mechanism or only the FFNs. The results (Table 4) show that both components contribute, but their effects differ. Penalizing only attention has a moderate impact on FLOPS, as FFNs still dominate. Penalizing only FFNs yields a larger FLOPS reduction but can be more detrimental to performance, as FFNs are believed to store factual knowledge geva2021transformer. The most effective strategy is penalizing both, allowing the model to find a flexible, optimal balance. This suggests the ”agents” in our system learn to coordinate their cost-saving strategies, contrasting with methods that apply uniform constraints, like fixed-rate pruning or hash layers singer2022hash.
@@ 5.3.3 Discussion: A New Perspective on Conditional Computation
The results offer a new perspective on conditional computation. While MoE models fedus2022switch achieve coarse-grained sparsity, our method induces a fine-grained, dynamic sparsity at the neuron and attention-weight level. It’s a continuous generalization of the discrete routing in MoE. Our framework also provides a principled way to control the performance-sparsity trade-off via the $\lambda$ parameter, a significant practical advantage over fixed architectural approaches. This work reframes interpretability research; instead of just observing what models do raghu2021vision, we can influence their behavior predictably. This proactive, interventionist approach, grounded in mechanism design narayanan2020learning, opens possibilities for building AI systems that are not only powerful but also efficient and transparent, with principles applicable to diverse fields from healthcare sensing 10804189 to secure federated learning .

### v3

@@ 5.3.2 Ablation Study: The Source of Savings
To disentangle the sources of efficiency gains, we applied the incentive penalty to only the attention mechanism or only the FFNs. The results (Table 4) show that both components contribute, but their effects differ. Penalizing only attention has a moderate impact on FLOPS, as FFNs still dominate. Penalizing only FFNs yields a larger FLOPS reduction but can be more detrimental to performance, as FFNs are believed to store factual knowledge geva2021transformer. The most effective strategy is penalizing both, allowing the model to find a flexible, optimal balance. This suggests the ”agents” in our system learn to coordinate their cost-saving strategies, contrasting with methods that apply uniform constraints, like fixed-rate pruning or hash layers.
@@ 5.3.3 Discussion: A New Perspective on Conditional Computation
The results offer a new perspective on conditional computation. While MoE models fedus2022switch achieve coarse-grained sparsity, our method induces a fine-grained, dynamic sparsity at the neuron and attention-weight level. It’s a continuous generalization of the discrete routing in MoE. Our framework also provides a principled way to control the performance-sparsity trade-off via the $\lambda$ parameter, a significant practical advantage over fixed architectural approaches. This work reframes interpretability research; instead of just observing what models do raghu2021vision, we can influence their behavior predictably. This proactive, interventionist approach, grounded in mechanism design narayanan2020learning, opens possibilities for building AI systems that are not only powerful but also efficient and transparent, with principles applicable to diverse fields from healthcare sensing to secure federated learning .